I don’t understand how to determine a standard deviation visually without doing the math for it.

90 Chapter 3Describing ualu b. What is the interquartile range in the number of prey types? Use the method outlined in Section 3.2 to calculate the quartiles. c. Can you calculate the mean number of prey types in the diet? Explain. 16. Francis Galton (1894) presented the following data on the flight speeds of 3207 old homing pigeons traveling at least 90 miles. 800 600 400 200 500 700 900 1100 1300 1500 Speed (yards/min) a. What tyne of aranh in thi

e. Examine the graph and visually determine the approximate value of the standard devi- ation (to the nearest 50 yards per minute) Explain how you obtained your estimate.
Answer #1

Ans:

Range Rule of Thumb:

The Range Rule of Thumb says that the range is about four times the standard deviation. The standard deviation is another measure of spread in statistics.

So,

Standard deviation=Range/4

For above data:

Range=1500-0=1500

Standard deviation=1500/4=375